OpenStudy (mayankdevnani):

A body of mass 2 kg is thrown up vertically with a kinetic energy of 490 J.If the acceleration due to gravity is 9.8 m/s^2, the height at which the kinetic energy of the body becomes half of the original value is - a) 50 m b) 25 m c) 12.5 m d) 10 m

OpenStudy (mayankdevnani):

@Hero @hartnn @Chlorophyll @robtobey @sauravshakya @kropot72 @jiteshmeghwal9 @hba

OpenStudy (mayankdevnani):

@mathslover have a look:)

mathslover (mathslover):

First find the original value ... what you get?

OpenStudy (mayankdevnani):

original value????

mathslover (mathslover):

sorry i think it is given

OpenStudy (anonymous):

Potential energy of the body =mgh

OpenStudy (anonymous):

u know it?

OpenStudy (mayankdevnani):

ok!!

OpenStudy (mayankdevnani):

yaa

mathslover (mathslover):

490 J = KE (ORIGINAL)

OpenStudy (mayankdevnani):

yaa

mathslover (mathslover):

half of it will be 490 /2 J = 245 J , correct?

OpenStudy (mayankdevnani):

why you do the half the K.E @mathslover

mathslover (mathslover):

@mayankdevnani because we have to find the height when the new kinetic energy equals half of the original one. ...

OpenStudy (mayankdevnani):

@03225186213

mathslover (mathslover):

therefore the new kinetic energy must be equal to 245 Joules.

OpenStudy (mayankdevnani):

yaa

OpenStudy (mayankdevnani):

correct!!!

mathslover (mathslover):

therefore we have : 1/2 m v^2 = 245 J 1/2 m (u^2 + 2gh ) = 245 J 1/2 m (2gh) = 245 mgh = 245 J 2 kg * 9.8 * h = 245

mathslover (mathslover):

solve for h and you will get h = 12.5 m

OpenStudy (mayankdevnani):

1/2 m (u^2 + 2gh ) = 245 J ??????????????

mathslover (mathslover):

got it?

OpenStudy (mayankdevnani):

no!!!!

OpenStudy (mayankdevnani):

1/2 m (u^2 + 2gh ) = 245 J this line!!!

mathslover (mathslover):

since v^2 = u^2 + 2gh therefore..

OpenStudy (mayankdevnani):

ok

OpenStudy (mayankdevnani):

thnx.... i got it @mathslover

mathslover (mathslover):

good,...

OpenStudy (mayankdevnani):

let @03225186213 be answer.... plz wait

OpenStudy (anonymous):

1/2 mv^2=mgh 490/2=2 (-9.8)h h= -9.8 because it is moving against gravity

mathslover (mathslover):

@03225186213 same I thought BUT it is given in the quest. that g = 9.8 m/s^2

OpenStudy (mayankdevnani):

yaaa you are correct thnx @mathslover and @03225186213

OpenStudy (anonymous):

welcome

OpenStudy (mayankdevnani):

:)